# Improved Algorithms for Contextual Dynamic Pricing

### v1

@@ 1.1 Related Work
Dynamic pricing has been extensively studied for half a century (Littlewood,, 1972; Rothstein,, 1974), leading to rich research on both theoretical and empirical fronts. For comprehensive surveys on the topic, we refer the readers to Bitran and Caldentey, (2003); Den Boer, (2015). While earlier works assumed that the buyer’s valuations are i.i.d. (Kleinberg and Leighton,, 2003; Besbes and Zeevi,, 2009; Keskin and Zeevi,, 2014; Cesa-Bianchi et al.,, 2019), recent research has increasingly focused on feature-based (or contextual) pricing problems. In this scenario, product value and pricing strategy depend on covariates. Pioneering works considered valuations depending deterministically on the covariates. Linear valuations have been the most studied (Amin et al.,, 2014; Javanmard and Nazerzadeh,, 2019; Cohen et al.,, 2020; Liu et al.,, 2021), yet a few authors have also explored non-parametric valuations (Mao et al.,, 2018).
Recent works have extended these methods to random valuations, mainly assuming that valuations are given by a function of the covariate, distorted by an additive i.i.d. noise. As this poses more challenges, authors have mostly focused on the simplest case of linear valuation functions, under additional assumptions. Initial studies assumed knowledge of the noise distribution (Cohen et al.,, 2020; Javanmard and Nazerzadeh,, 2019; Wang et al.,, 2023). This assumption was later relaxed, albeit with additional regularity requirements on the cumulative distribution function (c.d.f.) of the noise and/or the reward function Fan et al., (2024); Luo et al., (2024). Closest to our work, Xu and Wang, (2022) achieves a regret bound of $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$ for linear valuations, while assuming only the boundedness of the noise. Other parametric models have been explored, with, for example, generalized linear regression models Shah et al., (2019), though they also require strong assumptions, including quadratic behavior of the reward function around each optimal price. Few works have considered non-deterministic valuations with non-parametric valuation functions. Among those, Chen and Gallego, (2021) consider Lipschitz-continuous valuation functions of $d$-dimensional covariates. They achieve a regret of order $\tilde{\mathcal{O}}(T^{\nicefrac{{d+2}}{{d+4}}})$, assuming again quadratic behaviour around optimal prices. We refer to Table 1 for a comprehensive comparison between different previous works, their assumptions and regret bounds.
@@
Linear $F$ is known $\tilde{\mathcal{O}}(T^{\nicefrac{{2}}{{3}}})$ (Cohen et al.,, 2020)
$F$ is known or parametric, and log-concave $\tilde{\mathcal{O}}(T^{\nicefrac{{1}}{{2}}})$ Javanmard and Nazerzadeh, (2019)
$F$ has $m$-th order derivatives $\tilde{\mathcal{O}}(T^{\nicefrac{{2m+1}}{{4m-1}}})$ (Fan et al.,, 2024)
$F^{\prime\prime}$ is bounded $\tilde{\mathcal{O}}(T^{\nicefrac{{2}}{{3}}})\vee\|\theta-\widehat{\theta}\|_{1}T$ (Luo et al.,, 2024)
$F$ is Lipschitz $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$ (Fan et al.,, 2024), $\tilde{\mathcal{O}}(T^{\nicefrac{{2}}{{3}}})$ [this work]
$\Omega(T^{\nicefrac{{2}}{{3}}})$ (Xu and Wang,, 2022)
Bounded noise $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$, (Xu and Wang,, 2022)
Non-parametric $\pi(x,\cdot)$ is quadratic around its maximum for all $x$, $F$ and $g$ are Lipschitz $\tilde{\mathcal{O}}(T^{\nicefrac{{d+2}}{{d+4}}})$ (Chen and Gallego,, 2021)
$\Omega(T^{\nicefrac{{d+2}}{{d+4}}})$ (Chen and Gallego,, 2021)
@@
To improve on previous results, we design algorithms that share information on the noise distribution across different contexts. This idea relates to methods used in cross-learning, a research direction stemming from online bandit problems with graph feedback (Mannor and Shamir,, 2011; Alon et al.,, 2015). In this framework, introduced by Balseiro et al., (2019) and further studied in Schneider and Zimmert, (2024), when choosing to take action $i$ in context $x_{t}$, the agent observes the reward $r_{i}(x_{t})$ along with rewards $r_{i}(x_{t}^{\prime})$ associated with other contexts $x_{t^{\prime}}$. Our algorithms leverage similar principles to learn information usable across different contexts. However, compared to the typical problems addressed by cross-learning methods (e.g., first-price auctions, sleeping bandits, multi-armed bandits with exogenous costs), the contextual dynamic problem is more complex due to the intricate dependence of the reward on the unknown valuation function.
@@ 2.2 Assumptions
This assumption is classical in contextual dynamic pricing problems. We underline that contexts do not need to be random. In particular, they can be chosen by an adaptive adversary, aware of the seller’s strategy, and based on past realizations of $(x_{s},p_{s},\xi_{s})_{s<t}$. Assumption 1 is milder than the i.i.d. context assumption appearing in Fan et al., (2024); Shah et al., (2019); Chen and Gallego, (2021).
@@
2 is weaker than most of the assumptions in related works. For example, Javanmard and Nazerzadeh, (2019) require both $F$ and $1-F$ to be log-concave. Fan et al., (2024) assume that $F$ has $m$-th derivative, and that $\delta-\nicefrac{{1-F(\delta)}}{{F^{\prime}(\delta)}}$ is greater than some positive constant for all $\delta$, achieving a regret of order $\tilde{\mathcal{O}}(T^{\nicefrac{{2m+1}}{{4m-1}}})$. In the case $m=1$, they propose a different algorithm, reaching a regret $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$. Luo et al., (2024) consider Lipschitz-continuous noise, under the additional assumption that, for every $x$, $p^{*}(x)\in\operatorname*{arg\,max}_{p}\pi(x,p)$ is unique, and that $F^{\prime\prime}$ is bounded. Chen and Gallego, (2021) assume quadratic behaviour around every maxima: for every $x$, $p^{*}(x)\in\operatorname*{arg\,max}_{p}\pi(x,p)$, $p^{*}(x)$ is unique, and for all $p$, $C(p^{*}(x)-p)^{2}\leq\pi(x,p^{*}(x))-\pi(x,p)\leq C^{\prime}(p^{*}(x)-p)^{2}$ for some constants $C,C^{\prime}$. The only work considering non-Lipschitz c.d.f. is Xu and Wang, (2022); however, they achieve a higher regret bound of $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$.
@@ 2.3 Information Sharing in Contextual Dynamic Pricing
Another approach, akin to that used in Chen and Gallego, (2021), involves partitioning the covariate space into bins and running independent algorithms for non-parametric bandits (such as CAB1 (Kleinberg,, 2004)) within each bin. Let us assume, for simplicity, contexts in $[0,1]$, and that we partition this segment into $K$ bins. Then, the discretization error is $\nicefrac{{1}}{{K}}$. Classical results show that the regret in one bin is $\tilde{\mathcal{O}}(T_{K}^{\nicefrac{{2}}{{3}}})$, where $T_{K}=\nicefrac{{T}}{{K}}$ is the number of rounds in each bin. Consequently, the regret is $\tilde{\mathcal{O}}(\nicefrac{{T}}{{K}}+K\times(\nicefrac{{T}}{{K}})^{\nicefrac{{2}}{{3}}})$, which is minimized for $K=T^{\nicefrac{{1}}{{4}}}$, resulting in a regret $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$.
@@ 3.1 Outline of the Algorithm
Surprisingly, even though this estimation approach for $g(x_{t})$ from binary feedback is extremely simple to perform, this method appears to have never been used in dynamic pricing. Previous approaches for estimating valuation functions in the linear model include the regularized maximum-likelihood estimator (Javanmard and Nazerzadeh,, 2019; Wang et al.,, 2023), which requires knowledge of the noise distribution. Another approach used in Luo et al., (2024) relies on the relation between estimating a linear valuation function from binary feedback and the classical linear classification problem. The authors propose recovering the linear parameters $\theta$ through logistic regression; however, they do not provide an explicit estimation rate for $\theta$. Liu et al., (2021) use the EXP-4 algorithm to aggregate policies corresponding to different values of $\theta$ and $F$, thus circumventing the necessity to estimate them. In a similar vein, in the non-parametric valuation model, Chen and Gallego, (2021) avoid the need to estimate $g(x_{t})$ by employing independent bandit algorithms for each (binned) value of $x_{t}$. Closer to our method, Fan et al., (2024) also set uniform prices to obtain unbiased estimates of valuations. Yet, their analysis does not rely on classic bandit techniques, but rather uses more intricate arguments and requires stronger assumptions on $F$ and on the distribution of the contexts.
@@
The Price Elimination subroutine relies on the previous remark to select a price increment. For each increment $\delta_{k}$, we build a confidence bound $[\textnormal{LCB}_{t}(\delta_{k}),\textnormal{UCB}_{t}(\delta_{k})]=[p_{t}\widehat{D}^{k}_{t}\pm B_{y}(2L_{\xi}\epsilon+\sqrt{\nicefrac{{2\log(1/\alpha)}}{{N_{t}^{k}}}})]$ for the reward of price $p_{t}=\widehat{g}(x_{t})+\delta_{k}$. Then, we use a successive elimination algorithm (Even-Dar et al.,, 2006; Perchet and Rigollet,, 2013) to select a good increment. More precisely, we consider increments $\delta_{k}$ such that $\textnormal{UCB}_{t}(\delta_{k})\geq\max_{l}\textnormal{LCB}_{t}(\delta_{l})$, and we choose among these increments the increment $\delta_{k_{t}}$ that has been selected the least frequently. By doing so, we ensure to only select potentially optimal prices and gradually eliminate sub-optimal increments.
@@ 4 Linear Valuation Functions
11:              $\textnormal{LCB}_{t}(k)\leftarrow\left(\widehat{g}_{t}+k\epsilon\right)(\widehat{D}_{t}^{k}+\sqrt{\frac{2\log(\nicefrac{{1}}{{\alpha}})}{N_{t}^{k}}}-2L_{\xi}\epsilon)$
@@
14:         Update $\widehat{D}_{t+1}^{k_{t}}\leftarrow\frac{N^{k_{t}}_{t}\widehat{D}_{t}^{k_{t}}+o_{t}}{N^{k_{t},s}_{t}+1}$, $N^{k_{t}}_{t+1}\leftarrow N^{k_{t}}_{t}+1$.
@@
[See Appendix A for the full proof] Using Claim 1, we see that it is enough to prove that the Valuation Approximation phase allows to estimate $g(x_{t})$ up to precision $\epsilon=(\nicefrac{{d^{2}\log(T)^{2}}}{{T}})^{\nicefrac{{1}}{{3}}}$ in at most $O(d^{\nicefrac{{2}}{{3}}}T^{\nicefrac{{2}}{{3}}}\log(T)^{\nicefrac{{2}}{{3}}})$ rounds.
@@
Classical result on ridge regression in bandit framework (Abbasi-Yadkori et al.,, 2011) show that on a large probability event, $\|\theta-\widehat{\theta}_{t}\|_{\textbf{V}_{t}}\leq\left(B_{y}\sqrt{d\log\left(\nicefrac{{1+B_{x}^{2}T}}{{\alpha}}\right)}+B_{\theta}\right)$, so $|\widehat{g}(x_{t})-g(x_{t})|\leq\epsilon$.
@@
Theorem 1 provides a regret bound of order $\tilde{O}(T^{\nicefrac{{2}}{{3}}})$, showing that VAPE for Linear Valuations is minimax optimal, possibly up to sub-logarithmic terms and to sub-linear dependence in the dimension. Indeed, it matches the $T^{\nicefrac{{2}}{{3}}}$ lower bound established in Xu and Wang, (2022) for linear valuation functions and Lipschitz-continuous demand functions. This result represents a clear improvement over the existing regret bounds for the same problem. Indeed, VAPE achieves the regret bound conjectured in Luo et al., (2024) while at the same time removing their regularity assumption on the revenue function. On the other hand, we improve on the regret rate $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$ achieved respectively in Xu and Wang, (2022) under assumptions slightly milder than ours, and in Fan et al., (2024) under stronger assumptions.
@@ 5 Non-Parametric Valuation Functions
11:              $\textnormal{LCB}_{t}(k)\leftarrow\left(\widehat{g}_{t}+k\epsilon\right)(\widehat{D}_{t}^{k}+\sqrt{\frac{2\log(\nicefrac{{1}}{{\alpha}})}{N_{t}^{k}}}-2L_{\xi}\epsilon)$
@@
14:         Update $\widehat{D}_{t+1}^{k_{t}}\leftarrow\frac{N^{k_{t}}_{t}\widehat{D}_{t}^{k_{t}}+o_{t}}{N^{k_{t},s}_{t}+1}$, $N^{k_{t}}_{t+1}\leftarrow N^{k_{t}}_{t}+1$.
@@
[See Appendix B for the full proof] Using Claim 1, we only need to show that the length of the Valuation Approximation phase is at most of order $T^{\nicefrac{{d+2\beta}}{{d+3\beta}}}\log(T)^{\nicefrac{{\beta}}{{d+3\beta}}}$ and that w.h.p., it allows estimating $g$ uniformly on a ball of radius $B_{x}$ with precision $\epsilon\!=\!(\nicefrac{{T}}{{\log(T)}})^{\nicefrac{{-\beta}}{{d+3\beta}}}$.
@@
Theorem 2 shows that the Algorithm Valuation Approximation – Price Elimination for non-parametric valuations enjoys a $\tilde{O}(T^{\nicefrac{{d+2\beta}}{{d+3\beta}}})$ regret bound when the noise c.d.f. is Lipschitz and the valuation function Hölder-continuous. This result is the first of its kind under such minimal assumptions. In particular, previous work by Chen and Gallego, (2021) assumes quadratic behavior around the optimal price for all values of $g(x)$ – a very strong assumption. However, this rate is higher than the $\tilde{O}(T^{\nicefrac{{d+\beta}}{{d+2\beta}}})$ rates that are usually encountered in $\beta$-Hölder non-parametric bandits Bubeck et al., (2011). Thus, the question of optimality of the VAPE algorithmic scheme in the non-parametric valuation problem remains open.

### v2

@@ 1.1 Related Work
Dynamic pricing has been extensively studied for half a century [19, 26], leading to rich research on both theoretical and empirical fronts. For comprehensive surveys on the topic, we refer the readers to [6, 12]. While earlier works assumed that the buyer’s valuations are i.i.d. [18, 5, 16, 9], recent research has increasingly focused on feature-based (or contextual) pricing problems. In this scenario, product value and pricing strategy depend on covariates. Pioneering works considered valuations depending deterministically on the covariates. Linear valuations have been the most studied [3, 15, 11, 20], yet a few authors have also explored non-parametric valuations [24].
Recent works have extended these methods to random valuations, mainly assuming that valuations are given by a function of the covariate, distorted by an additive i.i.d. noise. As this poses more challenges, authors have mostly focused on the simplest case of linear valuation functions, under additional assumptions. Initial studies assumed knowledge of the noise distribution [11, 15, 30]. This assumption was later relaxed, albeit with additional regularity requirements on the cumulative distribution function (c.d.f.) of the noise and/or the reward function [14, 22], and then again by [31], that achieves a regret bound of $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$ for linear valuations, while assuming only the boundedness of the noise. Closest to our work [14, 21] also focus on the case in which the only regularity required is the Lipschitzness of the CDF. Their approaches show some similarities with our work but still achieve suboptimal regret rates. A more detailed comparison between ours an their algorithms is presented later on in the paper. Other parametric models have been explored, with, for example, generalized linear regression models [28], though they also require strong assumptions, including quadratic behavior of the reward function around each optimal price. Few works have considered non-deterministic valuations with non-parametric valuation functions. Among those, [10] consider Lipschitz-continuous valuation functions of $d$-dimensional covariates. They achieve a regret of order $\tilde{\mathcal{O}}(T^{\nicefrac{{d+2}}{{d+4}}})$, assuming again quadratic behaviour around optimal prices. We refer to Table 1 for a comprehensive comparison between different previous works, their assumptions and regret bounds.
@@
Linear $F$ is known $\tilde{\mathcal{O}}(T^{\nicefrac{{2}}{{3}}})$ [11]
$F$ is known or parametric, and log-concave $\tilde{\mathcal{O}}(T^{\nicefrac{{1}}{{2}}})$ [15]
$F$ has $m$-th order derivatives $\tilde{\mathcal{O}}(T^{\nicefrac{{2m+1}}{{4m-1}}})$ [14]
$F^{\prime\prime}$ is bounded $\tilde{\mathcal{O}}(T^{\nicefrac{{2}}{{3}}})\vee\|\theta-\widehat{\theta}\|_{1}T$ [22]
$F$ is Lipschitz $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$ [14, 21], $\tilde{\mathcal{O}}(T^{\nicefrac{{2}}{{3}}})$ [this work]
$\Omega(T^{\nicefrac{{2}}{{3}}})$ [31]
Bounded noise $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$, [31]
Non-parametric $\pi(x,\cdot)$ is quadratic around its maximum for all $x$, $F$ and $g$ are Lipschitz $\tilde{\mathcal{O}}(T^{\nicefrac{{d+2}}{{d+4}}})$ [10]
$\Omega(T^{\nicefrac{{d+2}}{{d+4}}})$ [10]
@@
To improve on previous results, we design algorithms that share information on the noise distribution across different contexts. This idea relates to methods used in cross-learning, a research direction stemming from online bandit problems with graph feedback [23, 2]. In this framework, introduced by [4] and further studied in [27], when choosing to take action $i$ in context $x_{t}$, the agent observes the reward $r_{i}(x_{t})$ along with rewards $r_{i}(x_{t}^{\prime})$ associated with other contexts $x_{t^{\prime}}$. Our algorithms leverage similar principles to learn information usable across different contexts. However, compared to the typical problems addressed by cross-learning methods (e.g., first-price auctions, sleeping bandits, multi-armed bandits with exogenous costs), the contextual dynamic problem is more complex due to the intricate dependence of the reward on the unknown valuation function.
@@ 2.2 Assumptions
This assumption is classical in contextual dynamic pricing problems. We underline that contexts do not need to be random. In particular, they can be chosen by an adaptive adversary, aware of the seller’s strategy, and based on past realizations of $(x_{s},p_{s},\xi_{s})_{s<t}$. Assumption 1 is milder than the i.i.d. context assumption appearing in [14, 28, 10].
@@
2 is weaker than most of the assumptions in related works. For example, [15] require both $F$ and $1-F$ to be log-concave. [14] assume that $F$ has $m$-th derivative, and that $\delta-\nicefrac{{1-F(\delta)}}{{F^{\prime}(\delta)}}$ is greater than some positive constant for all $\delta$, achieving a regret of order $\tilde{\mathcal{O}}(T^{\nicefrac{{2m+1}}{{4m-1}}})$. In the case $m=1$, they propose a different algorithm, reaching a regret $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$. [22] consider Lipschitz-continuous noise, under the additional assumption that, for every $x$, $p^{*}(x)\in\operatorname*{arg\,max}_{p}\pi(x,p)$ is unique, and that $F^{\prime\prime}$ is bounded. [10] assume quadratic behaviour around every maxima: for every $x$, $p^{*}(x)\in\operatorname*{arg\,max}_{p}\pi(x,p)$, $p^{*}(x)$ is unique, and for all $p$, $C(p^{*}(x)-p)^{2}\leq\pi(x,p^{*}(x))-\pi(x,p)\leq C^{\prime}(p^{*}(x)-p)^{2}$ for some constants $C,C^{\prime}$. The only work considering non-Lipschitz c.d.f. is [31]; however, they achieve a higher regret bound of $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$.
@@ 2.3 Information Sharing in Contextual Dynamic Pricing
Another approach, akin to that used in [10], involves partitioning the covariate space into bins and running independent algorithms for non-parametric bandits (such as CAB1 [17]) within each bin. Let us assume, for simplicity, contexts in $[0,1]$, and that we partition this segment into $K$ bins. Then, the discretization error is $\nicefrac{{1}}{{K}}$. Classical results show that the regret in one bin is $\tilde{\mathcal{O}}(T_{K}^{\nicefrac{{2}}{{3}}})$, where $T_{K}=\nicefrac{{T}}{{K}}$ is the number of rounds in each bin. Consequently, the regret is $\tilde{\mathcal{O}}(\nicefrac{{T}}{{K}}+K\times(\nicefrac{{T}}{{K}})^{\nicefrac{{2}}{{3}}})$, which is minimized for $K=T^{\nicefrac{{1}}{{4}}}$, resulting in a regret $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$.
@@ 3.1 Outline of the Algorithm
Previous approaches for estimating valuation functions in the linear model include the regularized maximum-likelihood estimator [15, 30], which requires knowledge of the noise distribution. Another approach used in [22] relies on the relation between estimating a linear valuation function from binary feedback and the classical linear classification problem. The authors propose recovering the linear parameters $\theta$ through logistic regression; however, they do not provide an explicit estimation rate for $\theta$. [20] use the EXP-4 algorithm to aggregate policies corresponding to different values of $\theta$ and $F$, thus circumventing the necessity to estimate them. In a similar vein, in the non-parametric valuation model, [10] avoid the need to estimate $g(x_{t})$ by employing independent bandit algorithms for each (binned) value of $x_{t}$. Closer to our method are the works of [14] and [21], who also set uniform prices to obtain unbiased estimates of the valuations. Nonetheless, their algorithms are significantly different from ours. First, they propose two-phased algorithms for which the phase length is set beforehand. Such an approach necessitates additional assumptions on how contexts are drawn; specifically, contexts are assumed to be i.i.d. from a distribution with a lower bound on the eigenvalues of the covariance matrix. This is needed to ensure that contexts observed in the first phase can represent the context distribution well. By contrast, our phases are adaptive, allowing our algorithm also to handle adversarial contexts and render these assumptions superfluous. Second, we obtain better regret rates by using piecewise-constant estimators, fitted in a regret-minimization sub-routine, as detailed in the next paragraph. On the other hand, [14] performs a phase of pure exploitation, relying on an estimate of the CDF $F$ that is constructed using Kernel methods. [21], instead, re-frames the problem as a perturbed linear bandit, which exhibits a regret linear in the dimension. However, this dimension depends on the size of the discretization grid – which is horizon dependent – leading to worse rates.
@@
The Price Elimination subroutine relies on the previous remark to select a price increment. For each increment $\delta_{k}$, we build a confidence bound $[\textnormal{LCB}_{t}(\delta_{k}),\textnormal{UCB}_{t}(\delta_{k})]=[p_{t}\widehat{D}^{k}_{t}\pm B_{y}(2L_{\xi}\epsilon+\sqrt{\nicefrac{{2\log(1/\alpha)}}{{N_{t}^{k}}}})]$ for the reward of price $p_{t}=\widehat{g}(x_{t})+\delta_{k}$. Then, we use a successive elimination algorithm [13, 25] to select a good increment. More precisely, we consider increments $\delta_{k}$ such that $\textnormal{UCB}_{t}(\delta_{k})\geq\max_{l}\textnormal{LCB}_{t}(\delta_{l})$, and we choose among these increments the increment $\delta_{k_{t}}$ that has been selected the least frequently. By doing so, we ensure to only select potentially optimal prices and gradually eliminate sub-optimal increments.
@@ 4 Linear Valuation Functions
11:              $\textnormal{LCB}_{t}(k)\leftarrow\left(\widehat{g}_{t}+k\epsilon\right)(\widehat{D}_{t}^{k}-\sqrt{\frac{2\log(\nicefrac{{1}}{{\alpha}})}{N_{t}^{k}}}-2L_{\xi}\epsilon)$
@@
14:         Update $\widehat{D}_{t+1}^{k_{t}}\leftarrow\frac{N^{k_{t}}_{t}\widehat{D}_{t}^{k_{t}}+o_{t}}{N^{k_{t}}_{t}+1}$, $N^{k_{t}}_{t+1}\leftarrow N^{k_{t}}_{t}+1$.
@@
[See Appendix B for the full proof] Using Claim 1, we see that it is enough to prove that the Valuation Approximation phase allows to estimate $g(x_{t})$ up to precision $\epsilon=(\nicefrac{{d^{2}\log(T)^{2}}}{{T}})^{\nicefrac{{1}}{{3}}}$ in at most $O(d^{\nicefrac{{2}}{{3}}}T^{\nicefrac{{2}}{{3}}}\log(T)^{\nicefrac{{2}}{{3}}})$ rounds.
@@
Classical result on ridge regression in bandit framework [1] show that on a large probability event, $\|\theta-\widehat{\theta}_{t}\|_{\textbf{V}_{t}}\leq\left(B_{y}\sqrt{d\log\left(\nicefrac{{1+B_{x}^{2}T}}{{\alpha}}\right)}+B_{\theta}\right)$, so $|\widehat{g}(x_{t})-g(x_{t})|\leq\epsilon$.
@@
Theorem 1 provides a regret bound of order $\tilde{O}(T^{\nicefrac{{2}}{{3}}})$, showing that VAPE for Linear Valuations is minimax optimal, possibly up to sub-logarithmic terms and to sub-linear dependence in the dimension. Indeed, it matches the $T^{\nicefrac{{2}}{{3}}}$ lower bound established in [31] for linear valuation functions and Lipschitz-continuous demand functions. This result represents a clear improvement over the existing regret bounds for the same problem. Indeed, VAPE achieves the regret bound conjectured in [22] while at the same time removing their regularity assumption on the revenue function. On the other hand, we improve on the regret rate $\tilde{\mathcal{O}}(T^{\nicefrac{{3}}{{4}}})$ achieved respectively in [31] under assumptions slightly milder than ours, and in [14] under stronger assumptions.
@@ 5 Non-Parametric Valuation Functions
11:              $\textnormal{LCB}_{t}(k)\leftarrow\left(\widehat{g}_{t}+k\epsilon\right)(\widehat{D}_{t}^{k}-\sqrt{\frac{2\log(\nicefrac{{1}}{{\alpha}})}{N_{t}^{k}}}-2L_{\xi}\epsilon)$
@@
14:         Update $\widehat{D}_{t+1}^{k_{t}}\leftarrow\frac{N^{k_{t}}_{t}\widehat{D}_{t}^{k_{t}}+o_{t}}{N^{k_{t}}_{t}+1}$, $N^{k_{t}}_{t+1}\leftarrow N^{k_{t}}_{t}+1$.
@@
[See Appendix C for the full proof] Using Claim 1, we only need to show that the length of the Valuation Approximation phase is at most of order $T^{\nicefrac{{d+2\beta}}{{d+3\beta}}}\log(T)^{\nicefrac{{\beta}}{{d+3\beta}}}$ and that w.h.p., it allows estimating $g$ uniformly on a ball of radius $B_{x}$ with precision $\epsilon\!=\!(\nicefrac{{T}}{{\log(T)}})^{\nicefrac{{-\beta}}{{d+3\beta}}}$.
@@
Theorem 2 shows that the Algorithm Valuation Approximation – Price Elimination for non-parametric valuations enjoys a $\tilde{O}(T^{\nicefrac{{d+2\beta}}{{d+3\beta}}})$ regret bound when the noise c.d.f. is Lipschitz and the valuation function Hölder-continuous. This result is the first of its kind under such minimal assumptions. In particular, previous work by [10] assumes quadratic behavior around the optimal price for all values of $g(x)$ – a very strong assumption. However, this rate is higher than the $\tilde{O}(T^{\nicefrac{{d+\beta}}{{d+2\beta}}})$ rates that are usually encountered in $\beta$-Hölder non-parametric bandits [7]. Thus, the question of optimality of the VAPE algorithmic scheme in the non-parametric valuation problem remains open.